OpenStudy (anonymous):

how to find the domain and range of f(x)=ln(x+sqrt x^2+1 ) @undeadknight26

OpenStudy (zzr0ck3r):

is it \(\log(x+\sqrt{x^2}+1)\)?

OpenStudy (zzr0ck3r):

if so then \(x+\sqrt{x^2}+1>0\) which is always true. Why?

hartnn (hartnn):

unlikely that would be the question.... \(\log(x+\sqrt{x^2+1})\) or most likely \(\log(x)+\sqrt{x^2+1}\)

OpenStudy (zzr0ck3r):

press enter already

OpenStudy (zzr0ck3r):

:P

hartnn (hartnn):

Tab + Enter

OpenStudy (anonymous):

but i cant understand on this math and here question

OpenStudy (zzr0ck3r):

what does tab + enter do?

hartnn (hartnn):

it posts

OpenStudy (zzr0ck3r):

o lol I see, I am tab+space bar

hartnn (hartnn):

that F, my dear, is not an odd function!

OpenStudy (zzr0ck3r):

we still need to know where that second bracket is.

hartnn (hartnn):

none of the 3 Fs we discussed are odd functions

OpenStudy (anonymous):

ummm oh but did you get understand this q or no ?

OpenStudy (zzr0ck3r):

no we don't know what the question is because you only have one (, we need you to close it ) example f(x) = 3x+2 is completely different than 3(x+2) so until you add that second parentheses we don't know what you are asking. we can only assume, and that is not safe in maths

hartnn (hartnn):

nopes, the question is not clear

OpenStudy (anonymous):

its not ( on last and its 1

OpenStudy (anonymous):

i edited on here (: ,,

OpenStudy (zzr0ck3r):

still not clear if you mean \(\sqrt{x^2+1}\) or \(\sqrt{x^2}+1\)

hartnn (hartnn):

i am quite sure its the 1st only then , F will be odd function

OpenStudy (anonymous):

okay.. do you know calculus ?

hartnn (hartnn):

yes

OpenStudy (anonymous):